Shift summary for Day shift 9/15/99

Julie Whitmore (jaws@fnal.gov)
Wed, 15 Sep 1999 15:36:41 -0500 (CDT)

Shift Summary for DAY shift, 15/Sept/99

On Shift: Ugur, Julie

Also present: Ed, BobT, Sydney, JasonH, Arun, SashaL, Valeri, Patty, JasonL,
Hogan, RickF, RickC, Dave Pushka, Doug, PeterSh, BudK, BrianL, Leo.

Hours of beam available at all intensities: 0.0
Hours of beam available at nominal intensity: 0.0
Hours of E832 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:none

========================================================================
Runs ended this shift:
none

Tapes brought to FCC this shift:
(including tapes for online split)

QKB152,QKK432-434,QKS432-434,QKC432-434
QKMU14-15,QKN039,QKNL36,QKP15-16,QKSM18-20,QKA027-29,
QKCC20-21,QKCR08,QKLA17-18,QKLS03

========================================================================
Accesses during shift:
Open access.

========================================================================
Overview of shift:
As found survey of detector in progress.
Lots of people work on switch-over to E799 configuration.

========================================================================
Description of Shift:

time comment
------------------------------------------------------------------------
7:05 Change pipeline software to E799 mode.

7:30 Survey crew shows up. Hogan LOTOs DCs. Open access starts.

8:00 Vacuum crew LOTOs gate valve for collimator survey.
AC maintenance people shut off AC to hall. Shut off Et.
Open highbay door to keep downstairs cool. Sasha and Valeri LOTO
TRDs and VV'. Bud/Brian deflate all helium bags except bag1.

8:05 Onsplit requests tape on ktev3 - QKPE16.

8:30 Pushka tries to fix MF15 interlock. Recalibrates vacuum reading
on G-P. This might not fix it.

8:50 Leonard tells us that the AC work will take the full day. BobT
takes down CsI RF system. He also starts work to swap out
CsI LV system Sector 10 -6.5V supply and Sector 8 -5V supply.

9:00 Vacuum is off.

9:35 Theo raises set point for KTeV Hall/Et temperature alarm to 82F.
He needs to reset it after AC work is complete.

10:00 Turn off HV to CsI/CA. Ron works on improving air seal on east side
of the array in an attempt to stabilize the temperature of the
array. Restore HV.

10:20 Doug LOTOs gate valve.

11:00 RC and Arun zero REG and ZCAL. Turn on MF15 to restore accidental
counter HVs. Doug informs us that the PC which runs the vacuum
system went down at ~7:15am. He resurrects it.

11:30 E799 collimator installed.

12:30 RickK coordinates software changes: ktevmon v4_01 is new public
version, new KQ bug fix for ktrig, change mmg hbook in ALR799.
Remake all code with ktevana v5_01.
Trigger changes: group 10 & 12 sources plugged in, Et thresholds
measured (not yet changed)..

15:20 Etta and Sydney take down ktev online cluster ktev1,2 and 4
for disk installation.

13:30 Brian and BudK install mylar window on DC4. Page Sasha and Valeri.
TRDs are rolled in and locked down.

14:30 Survey crew status: Finished DC1-4 "as found."
To do: Survey TRDs and realign DC4 (align x/y if it has moved
and take out the 64mil rotation about z that was seen in
"as found"). Finish "as found" for new collimator. Do "as found"
for regenerator. TRD and regenerator survey will continue
tomorrow morning.

14:45 BobT finishes work on CsI LV. There are currently no spares!

15:15 Leonard tells us that the AC work is done and AC is back on.

15:20 Etta and Sydney take down all of ktev online cluster for
disk partitioning.

15:25 Angela finishes taking DC gas samples.

========================================================================
Follow up Notes on Previous Shift

Finish flushing onsplit files for E832.

========================================================================
Items to Follow up on Next Shift

Keep an eye on KTeV hall temperature. Ashkan will reset temperature alarm
limit on KTeV Hall/Et temp serious alarm system after ktev cluster is back up.

Download RickK's 3MU3 information to RickT:
RickT needs to make 3MU3 hardline to L1 source patch panel connection and set
delay. See Log XXI p. 191.

Get BobT to turn on CsI LV and RF system.

Leave Et off for tomorrow's chilled water work.

Finish E832/E799 switch-over cake.